OpenStudy (anonymous):

What's the square root of 0.04?

Parth (parthkohli):

What number do you multiply by itself to get 0.04?

OpenStudy (anonymous):

2.

Parth (parthkohli):

Hint: you multiply 2 by itself to get 4.

Parth (parthkohli):

Hmm, but here, we have 0.04.

OpenStudy (anonymous):

I know

OpenStudy (anonymous):

Okay

OpenStudy (yanasidlinskiy):

@Elena_hayden What do you think it is??

OpenStudy (anonymous):

Idk 0.02???????

Parth (parthkohli):

\[\sqrt{0.04 }= \sqrt{\dfrac{4}{100}} = \dfrac{\sqrt{4}}{\sqrt{100}}\]

OpenStudy (yanasidlinskiy):

0.2 x 0.2 = 0.04

OpenStudy (anonymous):

I'm really bad with square roots

Parth (parthkohli):

What is the square root of 100?

OpenStudy (yanasidlinskiy):

Its ok. I think everyone is. Thats why were here!!!

OpenStudy (anonymous):

Lol thanks girl!

Parth (parthkohli):

lol

OpenStudy (yanasidlinskiy):

your welcome:)

OpenStudy (phi):

this is the same problem as the previous. write your decimal as something over 100 (se Parth's post) then take the square root of top and bottom

OpenStudy (anonymous):

.1414213562 was the answer I got using a calc